OpenStudy (lilsis76):

Determine Amplitude/period/phase shift /graph function. over 1 period. Indicate xintercepts and coordinates of Hi/low points on graph. y = cos (2x - pi/3) +1 y = A ( Bx - C ) so Amplitude is 1 Period: 2pi/B--> 2pi/2 = Pi Okay is my phase shift right? Phaseshift = C/B pi ---- 3 --------------- = pi/6? 2

OpenStudy (anonymous):

looks good to me i never remember the formula, so i set \(2x-\frac{\pi}{3}=0\) and solve for \(x\)

OpenStudy (anonymous):

this also help me remember whether the graph has moved left or right. in this case you get \(x=\frac{\pi}{6}\) so it is shifted to the right

OpenStudy (lilsis76):

oh!, okay so since it was positive it goes right, right? I thought if if was negative it goes right?? i am still confused on how they shift.

OpenStudy (lilsis76):